### 29 Feb 2024 v1

# MIMDRAM: An End-to-End Processing-Using-DRAM System for High-Throughput, Energy-Efficient and Programmer-Transparent Multiple-Instruction Multiple-Data Processing

Abstract
Processing-using-DRAM (PUD) is a processing-in-memory (PIM) approach that uses a DRAM array’s massive internal parallelism to execute very-wide (e.g., 16,384–262,144-bit-wide) data-parallel operations, in a single-instruction multiple-data (SIMD) fashion. However, DRAM rows’ large and rigid granularity limit the effectiveness and applicability of PUD in three ways. First, since applications have varying degrees of SIMD parallelism (which is often smaller than the DRAM row granularity), PUD execution often leads to underutilization, throughput loss, and energy waste. Second, due to the high area cost of implementing interconnects that connect columns in a wide DRAM row, most PUD architectures are limited to the execution of parallel map operations, where a single operation is performed over equally-sized input and output arrays. Third, the need to feed the wide DRAM row with tens of thousands of data elements combined with the lack of adequate compiler support for PUD systems create a programmability barrier, since programmers need to manually extract SIMD parallelism from an application and map computation to the PUD hardware.
Our goal is to design a flexible PUD system that overcomes the limitations caused by the large and rigid granularity of PUD. To this end, we propose MIMDRAM, a hardware/software co-designed PUD system that introduces new mechanisms to allocate and control only the necessary resources for a given PUD operation. The key idea of MIMDRAM is to leverage fine-grained DRAM (i.e., the ability to independently access smaller segments of a large DRAM row) for PUD computation. disable,disable,todo: disable,#D1MIMDRAM exploits this key idea to enable a multiple-instruction multiple-data (MIMD) execution model in each DRAM subarray (and SIMD execution within each DRAM row segment).
We evaluate MIMDRAM using twelve real-world applications and 495 multi-programmed application mixes. Our evaluation shows that MIMDRAM provides 34 the performance, 14.3 the energy efficiency, 1.7 the throughput, and 1.3 the fairness of a state-of-the-art PUD framework, along with 30.6 and 6.8 the energy efficiency of a high-end CPU and GPU, respectively. MIMDRAM adds small area cost to a DRAM chip (1.11%) and CPU die (0.6%). We hope and believe that MIMDRAM’s ideas and results will help to enable more efficient and easy-to-program PUD systems. To this end, we open source MIMDRAM at https://github.com/CMU-SAFARI/MIMDRAM.
1 Introduction
Data movement between computation units (e.g., CPUs, GPUs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and is expected to worsen due to the increasing data intensiveness of modern applications [21, 22]. To mitigate the overheads caused by data movement, several works propose processing-in-memory (PIM) architectures [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117]. There are two main approaches to PIM [8, 9]: (i) processing-near-memory (PNM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 104], where computation logic is added near the memory arrays (e.g., in a DRAM chip or at the logic layer of a 3D-stacked memory [118, 119, 120]); and (ii) processing-using-memory (PUM) [59, 60, 61, 62, 63, 64, 65, 66, 67, 121, 68, 69, 70, 71, 72, 52, 73, 74, 75, 76, 77, 78, 79, 122, 123, 80, 81, 82, 83, 124, 84, 85, 86, 87, 125, 88, 126, 127, 128, 89, 90, 129, 91, 92, 93, 94, 95, 130, 131, 132, 133, 134, 135, 136, 137, 138, 96, 139, 140, 97, 141, 142, 143, 98, 99, 144, 100, 101, 145, 102, 103, 105, 146], where computation is performed by exploiting the analog operational properties of the memory circuitry. There are two main advantages of PUM over PNM. First, PUM fundamentally eliminates data movement by performing computation in situ, while data movement still occurs between computation units and memory arrays in PNM. Second, PUM architectures exploit the large internal bandwidth and parallelism available inside the memory arrays, while PNM solutions are fundamentally bottlenecked by the memory’s internal data buses.
PUM architectures can be implemented using different memory technologies, including SRAM [71, 70, 96, 97, 98, 99], DRAM [125, 52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], emerging non-volatile [77, 60, 68, 82, 81, 90, 59] or NAND flash [80, 91, 92, 93, 94, 95]. Processing-using-DRAM (PUD) [52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], in particular, enables the execution of different bulk operations in DRAM (i.e., PUD operations), such as (i) data copy and initialization [74, 64, 75, 100], (ii) bitwise Boolean operations [61, 69, 67, 52, 63], (iii) arithmetic operations [66, 69, 63, 125, 101, 88], and (iv) lookup table based operations [78, 79, 145, 102, 103].
PUD architectures commonly employ bit-serial computation [125, 52, 69, 74, 64, 76, 73, 65, 61, 75, 62, 67, 101], where they map each data element of a PUD operation to a DRAM column.111We provide a detailed background on DRAM organization in section 2.1. A PUD architecture that leverages bit-serial computation effectively enables a very-wide single-instruction multiple-data (SIMD) execution model in DRAM, with a large and rigid computation granularity. The large computation granularity stems from the fact that bit-serial computation turns each one of the many DRAM columns in a DRAM subarray into a computing engine. For example, there are 16,384–262,144 DRAM columns in a DDR4 DRAM [147] subarray, and each can execute a single PUD operation over multiple data elements stored on the DRAM columns. The rigid computation granularity stems from the fact that the granularity at which DRAM rows are accessed dictates the computation granularity of a PUD operation. In commodity DRAM chips, all DRAM row accesses happen at a fixed granularity: the DRAM access circuitry addresses all DRAM columns in a DRAM row simultaneously. As such, every PUD operation operates simultaneously on all the 16,384 to 262,144 data elements (one per DRAM column) a DDR4 DRAM subarray stores, for example.
We highlight three limitations of state-of-the-art PUD architectures that stem from the large and rigid granularity of PUD’s very-wide SIMD execution model. First, to sustain high SIMD utilization (i.e., the fraction of SIMD lanes executing a useful operation), each PUD operation needs to operate on a large amount of data. Unfortunately, not all applications have the required large amount of data parallelism to sustain high SIMD utilization in PUD architectures. Our analysis of twelve general-purpose applications (compiled for a high-end CPU system) shows that the degree of SIMD parallelism an application has varies significantly, from as low as 8 to as high as 134,217,729 data elements per SIMD instruction (section 3). Second, the large granularity of PUD execution makes it costly to implement interconnects that connect columns in a DRAM row. Such interconnects could allow for the implementation of PUD operations that require shifting data across DRAM columns, which is a common computing pattern present in vector-to-scalar reduction operations, e.g., sum += A[i]. This limits PUD operations to only parallel map operations (i.e., operations where an output data of the same dimension as the input data is produced by applying a computation independently to each input operand). A prior work [63] proposes modifications to the DRAM array to enable communication across DRAM columns. However, the interconnection network that this work proposes can lead to prohibitive area cost in commodity DRAM chips (i.e., 21% DRAM area overhead, see section 8.3). Third, the need to feed a very-wide DRAM row with thousands of data elements combined with the lack of adequate compilation tools that can identify very-wide SIMD instructions in general-purpose applications and map such instructions to equivalent PUD operations create a programmability barrier for PUD architectures. In state-of-the-art PUD architectures [125, 52, 85, 66, 79, 69, 63, 88, 83, 86, 64, 65, 61, 62, 67, 101], the programmer needs to manually extract a large and fixed amount of data-parallelism from an application and map computation to the underlying PUD hardware, which is a daunting task.
Our goal is to design a flexible PUD substrate that overcomes the three limitations caused by the large and rigid granularity of PUD execution. To this end, we propose MIMDRAM, a hardware/software co-designed PUD system that introduces the ability to allocate and control only the required amount of computing resources inside the DRAM subarray for PUD computation. The key idea of MIMDRAM is to leverage fine-grained DRAM for PUD operations. Prior works on fine-grained DRAM [148, 149, 150, 151, 152, 153, 154, 155] leverage the hierarchical design of a DRAM subarray to enable DRAM row accesses with flexible granularity. A DRAM subarray is composed of multiple (e.g., 32–128) smaller 2D arrays of 512–1024 DRAM rows and 512–1024 columns, called DRAM mats [156, 62, 65, 150]. During a row access, the DRAM access circuitry simultaneously addresses all columns across all mats in a subarray, composing a large DRAM row of size [#columns_per_mat #mats_per_subarray]. Fine-grained DRAM modifies the DRAM access circuitry to enable reading/writing data from/to individual DRAM mats, allowing the access of DRAM rows with a smaller number of columns.
Inspired by fine-grained DRAM, we propose simple modifications to the DRAM access circuitry to enable addressing individual DRAM mats during PUD computation. Fine-grained DRAM for PUD computation brings four main advantages. First, by addressing individual DRAM mats, MIMDRAM better matches the granularity of a PUD operation to the available data-parallelism present in the application. Second, with fewer DRAM columns per PUD operation, MIMDRAM makes it feasible to implement low-cost interconnects inside a DRAM subarray, allowing for data movement within and across DRAM mats. Third, since the number of columns in a single DRAM mat is on par with the number of SIMD lanes in modern processors’ SIMD engines (e.g., 512 SIMD lanes in AVX-51 SIMD engines2 [157]), MIMDRAM can leverage traditional compilers to map SIMD instructions to PUD operations. Fourth, MIMDRAM leverages unused DRAM mats to concurrently execute independent PUD operations across DRAM mats in a single DRAM subarray. This enables the PUD substrate to execute a not-so-wide PUD operation in a subset of the DRAM mats and other independent PUD operations across the remaining DRAM mats within a single DRAM subarray, in a multiple-instruction multiple-data (MIMD) fashion [158, 159, 160].
MIMDRAM leverages fine-grained DRAM for PUD in hardware and software. On the hardware side, MIMDRAM proposes simple modifications to the DRAM subarray and includes new mechanisms to the memory controller that allow the execution of independent PUD operations across the DRAM mats in a single subarray. Concretely, MIMDRAM includes (i) latches, isolation transistors, and selection logic in the DRAM subarray’s access circuitry to enable the execution of independent PUD operations in different DRAM mats; (ii) two different low-cost interconnects placed respectively at the local and global DRAM I/O circuitry, which enable communication across columns of a DRAM row at varying granularities and the execution of vector-to-scalar reduction in DRAM at low hardware cost. In the memory controller, MIMDRAM includes a new control unit that orchestrates the concurrent execution of independent PUD operations in different mats of a DRAM subarray. On the software side, MIMDRAM implements compiler passes to (i) automatically vectorize code regions that can benefit from PUD execution (called PUD-friendly regions); (ii) for such regions, generate PUD operations with the most appropriate SIMD granularity; and (iii) schedule the concurrent execution of independent PUD operations in different DRAM mats. We discuss how to integrate MIMDRAM in a real system, including how MIMDRAM deals with (i) data allocation within a DRAM subarray and (ii) mapping of a PUD’s operands to guarantee high utilization of the PUD substrate.
We evaluate MIMDRAM’s performance, energy efficiency, throughput, fairness, and SIMD utilization for twelve real-world applications from four popular benchmark suites (i.e., Phoenix [161], Polybench [162], Rodinia [163], and SPEC 2017 [164]) and 495 multi-programmed application mixes. Our evaluation shows that MIMDRAM provides 34 the performance, 14.3 the energy efficiency, 1.7 the throughput, 1.3 the fairness, and 18.6 the SIMD utilization of SIMDRAM [101] (a state-of-the-art PUD framework); while providing 30.6 and 6.8 the energy efficiency of a high-end CPU and GPU, respectively. MIMDRAM’s improvements are due to its ability to (i) maximize the utilization of the PUD substrate by concurrently executing independent PUD operations within each DRAM subarray, and (ii) allocate only the necessary resources (i.e., appropriate number of DRAM mats) for a PUD operation. MIMDRAM adds small area cost to a state-of-the-art DRAM chip (1.11%) and state-of-the-art CPU die (0.6%).
We make the following key contributions:
To our knowledge, this is the first work to propose an end-to-end processing-using-DRAM (PUD) system for general-purpose applications, which executes operations in a multiple-instruction multiple-data (MIMD) fashion. MIMDRAM makes low-cost modifications to the DRAM subarray design that enable the concurrent execution of multiple independent PUD operations in a single DRAM subarray.
We propose compiler passes that take as input unmodified C/C++ applications and, transparently to the programmer, (i) identify loops that are suitable for PUD execution, (ii) transform the source code to use PUD operations, and (iii) schedule independent PUD operations for concurrent execution in each DRAM subarray, maximizing utilization of the underlying PUD architecture.
We evaluate MIMDRAM with a wide range of general-purpose applications and observe that it provides higher energy efficiency and system throughput than state-of-the-art PUD, CPU, and GPU architectures.
We open-source MIMDRAM at https://github.com/CMU-SAFARI/MIMDRAM.
2 Background
We first briefly explain the architecture of a typical DRAM chip.222We refer the reader to various prior works [16, 62, 14, 150, 61, 165, 64, 166, 167, 168, 169, 170, 155] for a more detailed description of the DRAM architecture. Second, we describe prior PUD works that MIMDRAM builds on top of.
2.1 DRAM Organization & Operation
DRAM Organization. A DRAM system comprises of a hierarchy of components, as Fig. 1 illustrates. A DRAM module (Fig. 1a) has several (e.g., 8–16) DRAM chips. A DRAM chip (Fig. 1b) has multiple DRAM banks (e.g., 8–16). A DRAM bank (Fig. 1c) has multiple (e.g., 64–128) 2D arrays of DRAM cells known as DRAM mats. Several DRAM mats (e.g., 8–16) are grouped in a DRAM subarray. In a DRAM bank, there are three global components that are used to access the DRAM mats (as Fig. 1c depicts): (i) a global row decoder that selects a row of DRAM cells across multiple mats in a subarray, (ii) a column select logic (CSL) that selects portions of the DRAM row based on the column address, and (iii) a global sense amplifier (i.e., global row buffer) that transfers the selected fraction of the data from the selected DRAM row through the global bitlines.
A DRAM mat (Fig. 1d) consists of a 2D array of DRAM cells organized into multiple rows (e.g., 512–1024) and multiple columns (e.g., 512–1024) [171, 172, 173]. A DRAM cell (Fig. 1e) consists of an access transistor and a storage capacitor. The source nodes of the access transistors of all the DRAM cells in the same column connect the cells’ storage capacitors to the same local bitline. The gate nodes of the access transistors of all the DRAM cells in the same row connect the cells’ access transistors to the same local wordline. Each mat contains (i) a local row decoder that drives the local wordlines to the appropriate voltage levels to open (activate) a row, (ii) a row of sense amplifiers (also called a local row buffer) that senses and latches data from the activated row, and (iii) helper flip-flops (HFFs) that drive a portion (e.g., ) of the data in the local row buffer to the global bitlines.
DRAM Operation. Three major steps are involved in serving a main memory request. First, to select a DRAM row, the memory controller issues an ACTIVATION (ACT) command with the row address. On receiving this command, the DRAM chip transfers all the data in the row to the corresponding local row buffer. Second, to access a cache line from the activated row, the memory controller issues a READ (RD) command with the column address of the request. Third, to enable the access of another DRAM row in the same bank, the memory controller issues a PRECHARGE (PRE) command with the address of the currently activated DRAM bank. This command disconnects the local bitline by disabling the local wordline, and the local bitline voltage is restored to its quiescent state.
2.2 Processing-Using-DRAM
In-DRAM-Row Copy. RowClone [64] enables copying a row to a row in the same subarray by issuing two consecutive ACT commands to these two rows, followed by a PRE command. This command sequence is called AAP [61].
In-DRAM AND/OR/NOT. Ambit [61, 62, 76, 74, 73, 65] shows that simultaneously activating three DRAM rows, via a DRAM operation called triple row activation (TRA), can perform in-DRAM bitwise AND and OR operations. Ambit defines a new command called AP that issues a TRA followed by a PRE. disable,disable,todo: disable,#A1Since TRA operations are destructive, Ambit divides DRAM rows into three groups for PUD computing: (i) the Data group, which contains regular data rows; (ii) the Control group, which consists of two rows (C0 and C1) with all-0 and all-1 values; and (iii) the Bitwise group, which contains six rows designated for computation (four regular rows, T0, T1, T2, T3; and two rows, DCC0 and DCC1, of dual-contact cells for NOT).
Generalizing In-DRAM Majority. SIMDRAM [101] proposes a three-step framework to implement PUD operations. In the first step, SIMDRAM converts an AND/OR/NOT-based representation of the desired operation into an equivalent optimized MAJ/NOT-based representation. disable,disable,todo: disable,#A1By doing so, SIMDRAM reduces the number of TRA operations required to implement the operation. In the second step, SIMDRAM generates the required sequence of DRAM commands to execute the desired operation. disable,disable,todo: disable,#A1Specifically, this step translates the MAJ/NOT-based implementation of the operation into AAPs/APs. This step involves (i) allocating the designated compute rows in DRAM to the operands and (ii) determining the optimized sequence of AAPs/APs that are required to perform the operation. This step’s output is a µProgram, i.e., the optimized sequence of AAPs/APs that will be used to execute the operation at runtime. Each µProgram corresponds to a different bbop instruction, which is one of the CPU ISA extensions to allow programs to interact with the SIMDRAM framework (see section 6.1). In the third step, SIMDRAM uses a control unit in the memory controller to execute the bbop instruction using the corresponding µProgram. SIMDRAM implements 16 bbop instructions, including abs, add, bitcount, div, max, min, mult, ReLU, sub, and-/or-/xor-reduction, equal, greater, greater equal, and if-else.
Fig. 2 illustrates how SIMDRAM executes a one-bit full addition operation using the sequence of row copy (AAP) and majority (AP) operations in DRAM. The figure shows one iteration of the full adder computation that computes Y = A + B + C. First, SIMDRAM uses a vertical data layout, where all bits of a data element are placed in a single DRAM column when performing PuD computation. Consequently, SIMDRAM employs a bulk bit-serial SIMD execution model, where each data element is mapped to a column of a DRAM row. This allows a DRAM subarray to operate as a PuD SIMD engine, where a single bit-serial operation is performed over a large number of independent data elements (i.e., as many data elements as the size of a logical DRAM row, for example, 65,536) at once. Second, as shown in the figure, each iteration of the full adder requires five AAPs ( , , , , ) and three APs ( , , ). A bit-serial addition of -bit operands needs iterations, thus APs and AAPs [101].
3 Motivation
The efficiency of state-of-the-art PUD substrates can be subpar when the SIMD parallelism that exists in an application is smaller than or not a multiple of the size of a DRAM row. To quantify the SIMD parallelism some real-world applications inherently possess, we profile the maximum vectorization factor of twelve real-world applications. The vectorization factor of a loop is the number of scalar operands that fit into a SIMD register [174, 175]. We calculate the maximum vectorization factor by multiplying the vectorization factor of a single loop iteration and the loop’s trip count [176]. We leverage modern compilers’ loop auto-vectorization engines, which allows us to have an initial understanding of the SIMD parallelism that a large number of applications possesses.
For our analysis, we use (i) the LLVM compiler toolchain [177] (version 12.0.0) to automatically vectorize loops in the application, and (ii) an LLVM pass [178, 179, 180] that instruments each application’s loop to, during execution, gather dynamic information about each vectorized loop, i.e., the loop trip count, execution count, execution time, and instruction breakdown [181]. We compile each application using the clang compiler [177], using the appropriate flags to enable the loop auto-vectorization engine and its loop vectorization report (i.e., -O3 -Rpass-analysis=loop-vectorize -Rpass=loop-vectorize).333See section 7 for the description of our applications and their dataset. We assume SIMDRAM [101] as the target PUD architecture.
Fig. 3 shows the distribution of maximum vectorization factors (y-axis) of all the vectorizable loops in an application (x-axis). disable,disable,todo: disable,#B3We indicate different amounts of SIMD parallelism with horizontal dashed lines for reference. We make two observations. First, the maximum vectorization factor varies both within an application and across different applications. Our analysis shows maximum vectorization factors as low as 8 and as high as 134,217,729. Second, only a small fraction of vectorized loops have enough maximum vectorization factor disable,disable,todo: disable,#B3(i.e., values above the green horizontal dashed line) to fully exploit the SIMD parallelism of SIMDRAM. On average, only 0.11% of all vectorized loops have a maximum vectorization factor equal to or greater than a DRAM row (i.e., greater than 65,536 data elements). We conclude that (i) real-world applications have varying degrees of SIMD parallelism; and (ii) these varying degrees of SIMD parallelism rarely take full advantage of the very-wide SIMD width of state-of-the-art PUD substrates.
Problem & Goal. We observe that the rigid granularity of PUD architectures limits their efficiency (and thus their effective applicability) for many applications. Such applications would benefit from a variable-size SIMD substrate that can dynamically adapt to the varying levels of SIMD parallelism (i.e., different vectorization factors) an application exhibits during its execution. Therefore, our goal is to design a flexible PUD substrate that (i) adapts to the varying levels of SIMD parallelism present in an application, and (ii) maximizes the utilization of the very-wide PUD engine disable,disable,todo: disable,#D1by concurrently exploiting parallelism across different PUD operations (potentially from different applications).
4 MIMDRAM: A MIMD PUD Architecture
MIMDRAM is a hardware/software co-designed PUD system that enables fine-grained PUD computation at low cost and low programming effort. The key idea of MIMDRAM is to leverage fine-grained DRAM activation for PUD, which provides three benefits. First, it enables MIMDRAM to allocate only the appropriate computation resources (based on the maximum vectorization factor of a loop) for a target loop, thereby reducing underutilization and energy waste. Second, MIMDRAM can currently execute multiple independent operations inside a single DRAM subarray independently in separate DRAM mats. This allows MIMDRAM to operate as a MIMD PUD substrate, increasing overall throughput. Third, MIMDRAM implements low-cost interconnects that enable moving data across DRAM columns across and within DRAM mats by combining fine-grained DRAM activation with simple modifications to the DRAM I/O circuitry. This enables MIMDRAM to implement reduction operations in DRAM without any intervention of the host CPU cores.
4.1 MIMDRAM: Hardware Overview
Fig. 4 shows an overview of the DRAM organization of MIMDRAM. Compared to the baseline Ambit subarray organization, MIMDRAM adds four new components (colored in green) to a DRAM subarray and DRAM bank, which enable (i) fine-grained PUD execution; (ii) global I/O data movement; and (iii) local I/O data movement.
Fine-Grained PUD Execution. To enable fine-grained PUD execution, MIMDRAM modifies Ambit’s subarray and the DRAM bank with three new hardware structures: the mat isolation transistor, the row decoder latch, and the mat selector. At a high level, the mat isolation transistor allows for the independent access and operation of DRAM mats within a subarray while the row decoder latch enables the execution of a PUD operation in a range of DRAM mats that the mat selector defines.
First, the mat isolation transistor ( in Fig. 4) segments the global wordline connected to the local row decoder in each DRAM mat in a subarray. Second, the row decoder latch ( ) stores the bits from the global wordline used to address the local row decoder. Third, the mat selector ( ), shared across all DRAM mats in a subarray, asserts one or more mat isolation transistors. The mat selector enables the connection between the global wordline and the row decoder latches belonging to a range of DRAM mats. When issuing PUD operations, the memory controller specifies the logical address of the first and last DRAM mats that the PUD operation targets (called logical mat range). Internally, each DRAM chip (i) identifies whether any of its DRAM mats belong to the logical mat range and (ii) translates the logical mat range into the appropriate physical mat range, which is used as input for the mat selector. With these structures, different PUD operations can execute in different ranges of DRAM mats.
For example, to execute a TRA in only , MIMDRAM performs four steps: (i) when issuing a TRA, the memory controller sends, alongside the row address information, the logical mat range [mat begin,mat end] = [#0,#0] to address ( in Fig. 4); (ii) the mat selector ( ) receives the logical mat range, translates it to the appropriate physical mat range, and raises the matline corresponding to , which asserts ’s mat isolation transistor ( ) and connects the global wordline to ’s row decoder latch; (iii) the bits of the global wordline used to drive ’s local row decoder are stored in ’s row decoder latch ( ); (iv) finally, ’s local row decoder drives the appropriate rows in ’s DRAM array based on the value stored in the row decoder latch. From here, the DRAM row activation (and thus, PUD computation) proceeds as described in section 2.1, only involving the DRAM rows in . Since the per-mat row decoder latch stores the local row address for a given row activation in a mat, the memory controller can issue a TRA to another DRAM mat while is being activated ( ).
Global I/O Data Movement. To enable data movement across different mats, MIMDRAM implements an inter-mat interconnect by slightly modifying the connection between the I/O bus and the global row buffer ( in Fig. 4). The inter-mat interconnect relies on the observation that the sense amplifiers in the global row buffer have higher drive than the sense amplifiers in the local row buffer [167, 182], allowing to directly drive data from the global row buffer into the local row buffer.444Prior work [182] leverages the same observation to copy DRAM columns from one subarray to another. To leverage this observation, MIMDRAM adds a 2:1 multiplexer to the input/output port of each set of four 1-bit sense amplifiers in the global row buffer. The multiplexer selects whether the data that is written to the sense amplifier set comes from the I/O bus or from the neighbor sense amplifier set .
To manage inter-mat data movement, MIMDRAM exposes a new DRAM command to the memory controller called GB-MOV (global I/O move). The GB-MOV command takes as input: (i) the logical mat range [mat begin,mat end], row address, and column address of the source DRAM row and column; and (ii) the logical mat range [mat begin,mat end], row address, and column address of the destination DRAM row and column. With the inter-mat interconnect and new DRAM command, MIMDRAM can move four bits555The number of bits the inter-mat interconnect can move at once depends on the number of HFFs already present in a DRAM mat. We assume that each mat has four HFFs, as prior works suggest [150, 151, 155]. of data from a source row and column (row, column) in to a destination row and column (row, column) in , in a DRAM subarray with DRAM mats. Once the memory controller receives a GB-MOV command, it performs three steps. First, the memory controller issues an ACT to the source row in , which loads the target DRAM row to ’s local sense amplifiers ( in Fig. 4). Concurrently, the memory controller issues an ACT to the destination row in , which connects row to ’s local sense amplifiers. Second, the memory controller issues a RD with the address of the four-bit source column to . The column select command loads the four-bit column from ’s local sense amplifiers to its HFFs, and ’s HFFs drive the corresponding set of four one-bit sense amplifiers in the global row buffer ( ). Third, the memory controller issues a WR with the address of the four-bit destination column to . Since the WR corresponds to a GB-MOV command, the multiplexer that connects ’s HFFs to the global row buffer takes as input the added datapath coming from instead of the conventional datapath coming from the I/O bus ( ). As a result, the data stored in is loaded into , which in turn drives ’s HFFs and local sense amplifiers ( ). Once the four-bit column coming from row is written into ’s local sense amplifiers, the local sense amplifiers finish the WR by restoring the local bitlines in to VDD or GND, thereby storing the four-bit column coming from column as a column of row ( ).
The conservative worst-case latency of a GB-MOV command (i.e., where the addresses of the source and the destination rows differ) is equal to ; where is latency from the start of row activation until the completion of the DRAM cell’s charge restoration, [182] is the latency of turning on the connection between the source and destination local sense amplifiers; is the minimum time interval between a WR and a PRE command, which allows the sense amplifiers to restore the data to the DRAM cells; is the latency between issuing a PRE and when the DRAM bank is ready for a new row activation.
Local I/O Data Movement. To enable data movement across columns within a DRAM mat, MIMDRAM implements an intra-mat interconnect ( in Fig. 4), which does not require any hardware modifications. Instead, it modifies the sequence of steps DRAM executes during a column access operation. There are two key observations that enable the intra-mat interconnect. First, we observe that the local bitlines of a DRAM mat already share an interconnection path via the HFFs and column select logic (as Fig. 5 illustrates). Second, the HFFs in a DRAM mat can latch and amplify the local row buffer’s data [167, 154].
To manage intra-mat data movement, MIMDRAM exposes a new DRAM command to the memory controller called LC-MOV (local I/O move). The LC-MOV command takes as input: (i) the logical mat range [mat begin,mat end] of the target row, (ii) the row and column addresses of the source DRAM row and column; and (iii) the row and column addresses of the destination DRAM row and column. With the intra-mat interconnect and new DRAM command, MIMDRAM can move four bits of data from a source row and column (row, column) to a destination row and column (row, column) in . Once the memory controller receives an LC-MOV command, it performs two steps, which Fig. 5 illustrates. In the first step, the memory controller performs an ACT–RD–PRE targeting row, column in . The ACT loads row to ’s local sense amplifier ( in Fig. 5). The RD moves four bits from row, as indexed by column, into the mat’s HFFs by enabling the appropriate transistors in the column select logic ( ). The HFFs are then enabled by transitioning the HFF enable signal from low to high. This allows the HFFs to latch and amplify the selected four-bit data column from the local sense amplifier ( ). The PRE closes row. Until here, the LC-MOV command operates exactly as a regular ACT–RD–PRE command sequence. However, differently from a regular ACT–RD–PRE, the LC-MOV command does not lower the HFF enable signal when the RD finishes. This allows the four-bit data from column to reside in the mat’s HFFs. In the second step, the memory controller performs an ACT–WR–PRE targeting row, column in . The ACT loads row into the mat’s local row buffer ( ), and the WR asserts the column select logic to column, creating a path between the HFFs and the local row buffer ( ). Since the HFF enable signal is kept high, the HFFs will not sense and latch the data from column. Instead, the HFFs overwrite the data stored in the local sense amplifier with the previously four-bit data latched from column. The new data stored in the mat’s local sense amplifier propagates through the local bitlines and is written to the destination DRAM cells ( ).
The conservative worst-case latency of an LC-MOV command (i.e., where the addresses of the source and the destination rows differ) is equal to .
4.1.1 PUD Vector Reduction
We describe how MIMDRAM uses the inter-mat and intra-mat interconnects to implement PUD vector reduction. To do so, we use a simple example, where MIMDRAM executes a vector addition followed by a vector reduction, i.e., out+=(A[i]+B[i]). We assume that DRAM has only two mats, and the data elements of the input arrays A and B are evenly distributed across the two DRAM mats, as Fig. 6 illustrates. MIMDRAM executes a vector reduction in three steps. In the first step, MIMDRAM executes a PUD addition operation over the data in the two DRAM mats ( ), storing the temporary output data C into the same mats where the computation takes place (i.e., C = {C[0], C[1]}). In the second step, MIMDRAM issues a GB-MOV to move part of the temporary output C[0] stored in to a temporary row tmp in (tmp C[0]) via the inter-mat interconnect ( – ), four bits (i.e., four data elements) at a time. MIMDRAM iteratively executes step 2 until all data elements of C[0] are copied to . In the third step, once the GB-MOV finishes, MIMDRAM executes the final addition operation, i.e. tmp + C[1], in . The final output of the vector reduction operation is stored in the destination row out in ( ).
Once the vector reduction operation finishes, the temporary output array stored in holds as many data elements as the number of DRAM columns in a mat (e.g., 512 data elements). MIMDRAM allows reducing the temporary output vector further to an output vector with four data elements using the intra-mat interconnect. The process is analogous to that employed during the 512-element vector reduction: MIMDRAM uses the intra-mat interconnect and the LC-MOV command to implement an adder tree inside a single DRAM mat.666The number of GB-MOV and LC-MOV commands issued depends on the bit-precision of the input operands [101].
4.2 MIMDRAM: Control & Execution
To enable MIMDRAM to execute in a MIMD fashion, we need to efficiently (i) encode and communicate information regarding the target DRAM mats (i.e., the target mat range) in a timely manner (i.e., respecting DRAM timing parameters) while (ii) orchestrating the execution of independent PUD operations across the DRAM mats of a DRAM subarray. To do so, we take a conservative design approach: we aim to integrate MIMDRAM in commodity DRAM chips by providing an implementation (i) compatible with existing DRAM standards and (ii) that does not add new pins to a DRAM chips.
Encoding MAT Information. MIMDRAM needs a compact way to encode the target mat information, since a DRAM module often contains many DRAM mats. To solve this issue, MIMDRAM only allows a PUD operation to be executed in a physically contiguous set of DRAM mats.777In section 6.3, we describe how we enforce physically contiguous mat allocation. In this way, when executing the DRAM commands (i.e., ACTs and PREs) that realize a PUD operation, the memory controller only needs to provide the first and last (logical) mats an ACT target. Then, MIMDRAM internally decides which (physical) mats fit into the provided mat range. To do so, MIMDRAM implements a simple chip select logic and mat identifier logic inside the I/O circuitry of each DRAM chip. The chip select logic and mat identifier logic take as input the logical mat range and output (i) if DRAM mats placed in a chip belong to the mat range, and (ii) the physical mat range. In case a DRAM mat placed in a chip belongs to the mat range, the DRAM chip queues the physical mat range in the mat queue (which we describe later in this section). The physical mat range is used as input for the mat selector (see Fig. 4). Since there are up to 128 DRAM mats in a DDR4 module [183], MIMDRAM uses 14 bits to encode the logical mat range (7 bits each for mat begin and mat end, each), from which (i) the three most significant bits are used to identify the target DRAM chip and (ii) the four least significant bits are used to identify individual mats. The chip select logic and mat identifier logic comprise simple hardware elements: four comparators, two 2-input AND gates, two 2:1 multiplexers, and a 3-bit chip id register in each DRAM chip.
Communicating MAT Information. MIMDRAM needs to communicate to the DRAM chip information regarding the target mats during a PUD operation. However, it is challenging to communicate the mat information alongside an ACT due to the narrow DRAM command/address (C/A) bus interface, since the memory controller uses most of the available pins during a row activation for row address and command communication.888There are 27 C/A pins in a DDR4 chip [184], from which only three pins are not used during an ACT command. Our key idea to solve this issue is to overlap the latency of communicating the mat information to DRAM with the latency of DRAM commands in a µProgram in two ways: (i) ACT–ACT overlap, and (ii) PRE–ACT overlap. The first case (ACT–ACT overlap) happens when issuing a row copy operation (AAP). In this case, the mat information required by the second ACT is transmitted immediately after issuing the first ACT, exploiting the delay between two activations. The mat information is buffered once it reaches DRAM. The second case (PRE–ACT overlap) happens when issuing the first ACT in a row copy operation or the ACT in a TRA. We notice that (i) the first ACT command in an AAP/AP is always preceded by a PRE (due to a previous AAP/AP, or due to a previous DRAM request), and (ii) a PRE does not use the row address pins, since it targets a DRAM bank (not a DRAM row). Thus, MIMDRAM uses the row address pins during a PRE that immediately precedes the first ACT in an AAP/AP command sequence to communicate the mat information.999disable,disable,todo: disable,#A2If there are insufficient pins in the DDRx interface to communicate mat information (e.g., as in DDR5 [185]), MIMDRAM utilizes multiple DRAM C/A cycles to propagate the mat information. For example, in DDR5, MIMDRAM still performs PRE-ACT overlap, communicating the mat information in two cycles. Note that an extra cycle does not impact MIMDRAM’s performance, since in a PRE-ACT command sequence, the PRE still needs to wait for the completion of the ACT for more than two DRAM C/A cycles.
Timing of MAT Information. MIMDRAM needs to communicate the mat information before a respective ACT in a µProgram. Communicating the mat information immediately after the memory controller issues the ACT would open the entire DRAM row (instead of only the relevant portion of the DRAM row). To solve this issue, we devise a simple queuing-based mechanism for partial row activation. Our mechanism relies on the fact that the execution order of ACTs and PREs in a µProgram is deterministic.101010To realize a PUD operation, the memory controller must respect the order in which ACT and PRE commands are specified in the µProgram. Therefore, during PUD execution, ACTs and PREs in a µProgram cannot be reordered, and the behavior of the µProgram is thus deterministic. If the memory controller is performing maintenance operation to a DRAM bank, the AAP/AP commands of a PUD operation wait until the maintenance operation finishes. Thus, we can add to each DRAM command in an AAP/AP the information about when the DRAM circuitry should propagate the mat information. MIMDRAM leverages this key idea by adding a mat queue to the I/O logic of each DRAM chip and adding extra functionality to the existing ACT and PRE commands to control the mat queue: (i) ACT-enqueue issues an ACT to row_addr in the first DRAM clock cycle and enqueues [mat_begin,mat_end] in the second DRAM clock cycle; (ii) PRE-enqueue issues a PRE to bank_id and enqueues [mat_begin,mat_end]; (iii) ACT-dequeue issues an ACT to row_addr and dequeues from the mat queue.
Orchestrating MAT Information. MIMDRAM needs to execute different PUD operations concurrently. To this end, we implement a control unit inside the memory controller on the CPU die, which Fig. 7 illustrates. MIMDRAM leverages SIMDRAM control unit to translate each bbop instruction into its corresponding µProgram and adds extra circuitry to (i) schedule each µProgram based on its target mats and (ii) maintain multiple µProgram contexts. MIMDRAM control unit consists of four main components. First, bbop buffer, which stores bbops dispatched by the host CPU. Second, mat scheduler, which schedules the most appropriate bbop to execute depending on the bbop’s mat range and current mat utilization. Third, mat scoreboard, which tracks whether a given mat is being used by a bbop instruction. The mat scoreboard stores an -bit mat bitmap that keeps track of which mats are currently in use, where is the number of mats in the DRAM module. The mat scoreboard can index a range of positions in the mat bitmap using a mat index. Fourth, several (e.g., eight) µProgram processing engines, each of which translates a bbop into its respective µProgram and controls the bbop’s execution.
MIMDRAM control unit works in four steps. In the first step, MIMDRAM control unit enqueues an incoming bbop instruction dispatched by the host CPU ( in Fig. 7) in the bbop buffer. In the second step, the mat scheduler scans the bbop buffer from the oldest to the newest element. Then, the mat scheduler employs an online first fit algorithm [186] to select a bbop to be executed. For each bbop in the bbop buffer, the algorithm: (i) extracts the mat range information encoded in the bbop ( ), which is used to index the mat scoreboard ( ); (ii) reads the mat bitmap to identify whether the mats belonging to the bbop’s mat range are currently free or busy ( ); (iii) in case the mats are free, the mat scheduler writes a new mat bitmap to the mat scoreboard, indicating that the given mat range is now busy, selects the current bbop to be executed by allocating and copying the bbop to a free µProgram processing engine ( ), and removes the current bbop from the bbop buffer ( ); (iv) in case the mats belonging to the bbop’s range are busy, the mat scheduler reads the next available bbop from the bbop buffer and repeats (i)–(iii). In the third step, one or multiple µProgram processing engines execute their allocated bbop, issuing AAPs/APs to the DRAM chips ( ). The µProgram processing engine is disable,disable,todo: disable,#A4responsible for maintaining the timing of AAP/AP commands. In our design, we avoid the need to maintain state for all DRAM mats in a DRAM module individually by: (i) only allowing a PUD operation to address a contiguous range of DRAM mats, which share state as they execute the same sequence of ACT-PRE commands and (ii) limiting the number of concurrent PUD operations to the number of µProgram processing engines available in the control unit. In the fourth step, when a µProgram processing engine finishes executing, it frees its allocated mats by correspondingly updating the mat bitmap in the mat scoreboard ( ) and notifies the CPU that the execution of the bbop instruction is done.
5 MIMDRAM: Software Support
To ease MIMDRAM’s programmability, we provide compiler support to transparently map SIMD operations to MIMDRAM. Fig. 8 illustrates MIMDRAM’s compilation flow, which we implement using LLVM [177]: we take a C/C++ application’s source code as input, perform three transformations passes, and output a binary with a mix of CPU and PUD instructions.
Pass 1: Code Identification. The first pass is responsible for code identification. Its goal is to identify (i) loops that can be successfully auto-vectorized and (ii) the appropriate vectorization factor of a given vectorized loop. The code identification pass takes as input the application’s LLVM intermediate representation (IR) generated by the compiler’s front-end. It produces as output an optimized IR containing SIMD instructions that will be translated to bbop instructions. We leverage the native LLVM’s loop auto-vectorization pass [187] to identify and transform loops into their vectorized form ( in Fig. 8).111111Prior works [188, 189] also leverage modern compilers’ loop auto-vectorization engines to generate instructions to processing-near-memory (PNM) architectures equipped with SIMD engines. We apply two modifications to LLVM’s loop auto-vectorization pass. First, instead of using a cost model to choose the vectorization factor that leads to the highest performance improvement compared to a scalar version of the same loop, we always select the maximum vectorization factor for the loop. disable,disable,todo: disable,#C2This is important because the native cost model takes into account the hardware characteristics of a target CPU SIMD engine (i.e., number of available vector registers, SIMD width of the target execution engine, the latency of different SIMD instructions), which are not representative of our MIMDRAM engine with a variable SIMD width. Second, we modify the code generation routine for a given vectorized loop. Concretely, for a given vectorized loop, we identify and remove memory instructions related to each arithmetic SIMD operation (i.e., load/store instructions that manipulate vector registers) since PUD operations directly manipulate the data stored in DRAM; disable,disable,todo: disable,#C2thus, there is no need to explicitly move data into/out SIMD registers.
Pass 2: Code Scheduling & Data Mapping. The second pass is responsible for code scheduling and data mapping. Its goal is to improve overall SIMD utilization by allowing the distribution of independent PUD instructions across DRAM mats. Since PUD instructions operate disable,disable,todo: disable,#C2directly on the data stored in DRAM, the DRAM mat where the data is allocated determines the efficiency and utilization of the PUD SIMD engine. If operands of independent instructions are distributed across different DRAM mats, such instructions can be executed concurrently. Likewise, operands of dependent instructions are mapped to the same DRAM mat. In that case, intermediate data that one instruction produces and the next instruction consumes do not need to be moved across different DRAM mats, improving energy efficiency. Leveraging these observations, the code scheduling pass takes as input all bbop instructions the code identification pass generates and outputs new bbop instructions containing metadata regarding their mat location (i.e., mat label). The code scheduling pass works in two steps.
In the first step, the code scheduling pass creates a data-dependency graph (DDG) of the vectorized instructions ( ). disable,disable,todo: disable,#C2Each node represents a bbop instruction, incoming edges represent input, and outgoing edges represent output of the bbop. In the second step, the code scheduling pass takes as input the DDG and employs a data scheduling algorithm to distribute bbop instructions across DRAM mats ( ). The data scheduling algorithm traverses the DDG in topological order to respect dependencies between bbop instructions using a depth-first search (DFS) kernel, which is a common algorithm for topological ordering [190, 191], and performs three operations. First, the algorithm traverses the left nodes in the DDG, assigning a single mat label i to nodes in the left path ( - ). Second, when the algorithm reaches a leaf node, it traverses the right sub-tree in the DDG. In this case, the algorithm assigns a new mat label j to the nodes in the right path in the sub-tree ( - ). Third, once the algorithm visits all the nodes in the right sub-tree, it returns to the parent node of the sub-tree. Since the parent node has already been visited when descending into the left path, the left and right sub-tree nodes will be assigned to different mats while having data dependencies across them (as indicated by the parent node). In this case, the algorithm creates a data movement bbop instruction (see section 6.1) to move the output produced by the right sub-tree from mat label j to mat label i ( - ). This process disable,disable,todo: disable,#C2repeats until the algorithm visits all nodes in the DDG.
Pass 3: Code Generation. The third pass is responsible for (i) data allocation and (ii) code generation. It takes as input the LLVM IR containing both CPU and bbop instructions (with metadata) and produces a binary to the target ISA ( ). To implement data allocation, the code generation pass first identifies calls for memory allocation routines (e.g., malloc) associated with operands of bbops and replaces such memory allocation routines with a specialized PIM memory allocation routine (i.e., pim_malloc, see section 6.3). pim_malloc receives as input the mat label assigned to its associated bbop instruction. Second, the pass inserts a bbop_trsp_init instruction right after each pim_malloc call for each memory object that is an input/output of a bbop instruction. This instruction registers the memory object in MIMDRAM’s transposition unit (section 6.2). Similar to the pim_malloc call, the bbop_trsp_init instruction receives as input the mat label assigned to its associated bbop instruction. To implement code generation, we modify LLVM’s X86 back-end to identify bbop instructions and generate the appropriate assembly code. In case the application uses parallel primitives (e.g., OpenMP pragmas [192]) to parallelize outermost loops, the code generation pass interacts with the underlying runtime system to statically distribute bbop instructions from innermost loops across the available DRAM mats in a subarray, i.e., mats with unassigned mat labels. disable,disable,todo: disable,#C2This allows MIMDRAM to execute in a single-instruction multiple-thread (SIMT) [193, 194] fashion for manually parallelized applications.
6 System Support for MIMDRAM
We envision MIMDRAM as a tightly-coupled accelerator for the host processor. As such, MIMDRAM relies on the host processor for its system integration, which includes ISA support (section 6.1), instruction execution & data transposition (section 6.2), and operating system support for address translation and data allocation & alignment (section 6.3).
6.1 Instruction-Set Architecture
Table 1 shows the CPU ISA extensions that MIMDRAM exposes to the compiler.121212disable,disable,todo: disable,#E3MIMDRAM ISA extensions are vector-oriented by design. We did not use an existing ISA because we needed to define new fields for MIMDRAM that do not exist in current vector ISAs (e.g., mat label information). disable,disable,todo: disable,#D6Instead, we propose to extend the baseline CPU ISA with MIMDRAM instructions since there is usually more than enough unused opcode space to support the extra opcodes that MIMDRAM requires [195, 196]. Extending the CPU ISA to interface with accelerators is a common approach [101, 18, 61, 197, 198]. There are five types of instructions: (i) object initialization instructions, (ii) 1-input arithmetic instructions, (iii) 2-input arithmetic instructions, (iv) predication instructions, and (v) data movement instructions. The first three types of MIMDRAM instructions are inherited from the SIMDRAM ISA [101]. These instructions can be further divided into two categories: (i) operations with one input operand (e.g., bitcount, ReLU), and (ii) operations with two input operands (e.g., addition, division, equal, maximum). To enable predication, MIMDRAM uses the bbop_if_else instruction that SIMDRAM introduces, which takes as input three operands: two input arrays (src and src) and one predicate array (select). We modify such instructions by including two new fields: (i) mat label (ML), which identifies groups of instructions that must execute inside the same DRAM mat, and (ii) vectorization factor (VF), which dictates how many scalar operands are packed within the vector instruction. These two new fields are automatically generated by MIMDRAM’s compiler passes (section 5).
| Type | ISA Format |
| Initialization | bbop_trsp_init addr, size, n, ML |
| 1-Input Arith. | bbop_op dst, src, size, n, ML, VF |
| 2-Input Arith. | bbop_op dst, src, src size, n, ML, VF |
| Predication | bbop_if_else dst, src, src, sel, size, n, ML, VF |
| Data Move | bbop_mov dst, dst_idx, src, src_idx, size, n |
Data movement instructions allow the compiler to trigger inter-mat and intra-mat data movement operations. In a data movement instruction, dst and src represent the source and destination arrays; dst_idx and src_idx represent the first position of the first element inside the source and destination arrays to be moved; size represents the number of elements to move from source to the destination array; n represents the number of bits in each array element. MIMDRAM control unit automatically identifies the mat range the data movement instruction targets by calculating the distance between the source and destination arrays, taking into account the indexes and number of elements to move. In case the source and destination mats are the same, MIMDRAM control unit translates the data movement instruction into an LC-MOV command: otherwise, a GB-MOV command.
6.2 Execution & Data Transposition
Instruction Fetch and Dispatch. MIMDRAM relies on the host CPU to offload bbop instructions to DRAM since they are part of the CPU ISA. Assuming that the host CPU consists of one or more out-of-order cores, MIMDRAM leverages the host processor’s front-end to (i) identify and (ii) dispatch to MIMDRAM control unit only independent bbops. This simplifies the design of MIMDRAM control unit since no in-flight bbop instructions will have data dependencies. As a result, MIMDRAM control unit can freely schedule µPrograms to the PUD SIMD engine as they arrive.
Data Coherence. Input arrays to MIMDRAM may be generated or modified by the CPU, and the data updates may reside only in the cache (e.g., because the updates have not yet been written back to DRAM). To ensure that MIMDRAM does not operate on stale data, programmers are responsible for flushing cache lines [199, 200] modified by the CPU. MIMDRAM can leverage coherence optimizations tailored to PIM to improve overall performance [38, 44].
MIMDRAM Transposition Unit. MIMDRAM transposition unit shares the same hardware components and functionalities as the SIMDRAM transposition unit [101], which includes: (i) object tracker, a small cache that keeps track of the memory objects used by bbop instructions; (ii) an horizontal to vertical transpose unit, which converts cache lines of memory objects stored in the object tracker from a horizontal to vertical data layout during a last-level cache (LLC) writeback; (iii) a vertical to horizontal transpose unit, which converts cache lines of memory objects stored in the object tracker from a vertical to horizontal data layout during an LLC read request; (iv) store and fetch units, which generate memory read/write requests using the transpose units’ output data. One main limitation of the SIMDRAM transposition unit is that it needs to fill at least an entire DRAM row with vertically-laid out data before the execution of a bbop. Instead, MIMDRAM transposes only as much data as required to fill the segment of the DRAM row that the bbop instruction operates over. To do so, the MIMDRAM transposition unit adds information regarding the mat range a memory object operates to the object tracker.
6.3 Operating System Support
Address Translation. As SIMDRAM, MIMDRAM operates directly on physical addresses. When the CPU issues a bbop instruction, the instruction’s virtual memory addresses are translated into their corresponding physical addresses using the same translation lookaside buffer (TLB) lookup mechanisms used by regular load/store operations.
Data Allocation & Alignment. disable,disable,todo: disable,#CQ1MIMDRAM (as other PUD architectures [64, 75, 79, 61, 62, 76, 74, 73, 65, 86]) requires OS support to guarantee that data is properly mapped and aligned within the boundaries of the bank/subarray/mat that will perform computation. Particularly, since PUD operations are executed in-situ, it is essential to enforce that memory objects belonging to the same bbop (and their dependent instructions) are placed together in the same DRAM mats. To achieve this functionality, we propose the implementation of a new data allocation API called pim_malloc. The main idea is to allow the compiler to inform the OS memory allocator about the memory objects that must be allocated inside the same set of DRAM mats. The pim_malloc API takes as input the size of the memory region to allocate (as a regular malloc instruction) and the mat label that the compiler generates (section 5). Then, it ensures that all memory objects with the same mat label are placed together within a set of DRAM mats that satisfies the target memory allocation size.
To allow the pim_malloc API to influence the OS memory allocator and ensure that memory objects are placed within specific DRAM mats, disable,disable,todo: disable,#CQ1we propose a new lazy data allocation routine (in the kernel) for pim_malloc objects. This routine has three main components: (i) information about the DRAM organization (e.g., row, column, and mat sizes), (ii) the DRAM interleaving scheme, which the memory controller provides via an open firmware device tree [201];131313The DRAM interleaving scheme can be obtained by reverse engineering the bit locations of memory addresses [202, 203, 204, 205]. disable,disable,todo: disable,#A3Even though typical DRAM interleaving does not take mats into account, it is relatively straightforward to reverse-engineer how a memory address is distributed across the DRAM mats in a DRAM module, since the mat interleaving is a function of the DRAM chip’s organization. For example, in a DDR4 module with 8 chips, 16 mats per chip, and 4 HFFs per mat, a 64 B cache line is evenly distributed across all 128 total mats; i.e., the four least-significant bits of the cache line are placed in mat 0, chip 0, and the four most-significant bits of the cache line are placed in mat 15, chip 7. Our pim_malloc API takes into account such mat interleaving. and (iii) a huge page pool for pim_malloc objects (configured during boot time), which guarantees that virtual addresses assigned to a pim_malloc object are contiguous in the physical address space and that DRAM mats are free whenever a pim_malloc object is allocated. The allocation routine uses the DRAM address mapping knowledge to split the huge pages into different memory regions. Then, when an application calls the pim_malloc API, the allocation routine selects the appropriate memory region that satisfies pim_malloc. disable,disable,todo: disable,#D7Internally, the pim_malloc API operates using three main sub-tasks, depending on the order of the data allocation: (i) pim_preallocate, for data pre-allocation; (ii) pim_alloc, for the first data allocation; and (iii) pim_alloc_align, for subsequent aligned allocations.
(i) Pre-Allocation. The first sub-task’s role is to indicate the number of huge pages available for PUD allocations. We leave it to the user to provide the number of huge pages used for PUD operations because huge pages are scarce in the system.
(ii) First Allocation. The second sub-task uses the worst-fit allocation scheme [206] to manage the allocation of memory regions in the huge page pool. The key idea behind this placement strategy is to optimize the remaining memory space after allocations to increase the chances of accommodating another process in the remaining space. For the first PUD memory allocation, the pim_alloc sub-task simply scans an ordered array data structure (similar to the one used in the Linux Kernel buddy allocator algorithm [207], where each entry represents the number of memory regions in a single subarray) to select the subarray with the largest amount of memory regions available. If the requested memory allocation requires more than one memory region, MIMDRAM iteratively scans the ordered array, searching for the next largest memory region until the memory allocation is fully satisfied. Once enough space is allocated, pim_alloc sub-task creates a new allocation object and inserts it in an allocation hashmap, which is indexed by the allocation’s virtual address. The sub-task needs to keep track of allocations since it might need to find a memory region from the same subarray/mat when performing future aligned allocations.
(iii) Aligned Allocation. After allocating a memory region for the first operand in a PUD operation, the user can use this memory region as a regular memory object. However, when allocating the remaining operands for a PUD operation, the pim_malloc API needs to guarantee data alignment for all memory objects within the same DRAM subarray/mat. To this end, the third sub-task (pim_malloc_align) identifies a previously allocated memory region to which the current memory allocation must be aligned (based on the compiler-generated mat labels). The pim_malloc_align sub-task works in five main steps. First, it searches the allocation hashmap for a match with previously allocated memory regions. If a match is not found, the allocation fails. Second, if a match is found, the pim_malloc_align sub-task iterates through the identified previously-allocated memory regions. Third, for each memory region, the sub-task identifies its source subarray/mat address and tries to allocate another memory region in the same subarray/mat for the new allocation. Fourth, if the subarray/mat of a given memory region has no free region, the sub-task allocates a new memory region from another subarray/mat following the worst-fit allocation scheme. Since we use a worst-fit allocation scheme that always selects the largest number of memory regions available during memory allocation for the first operand of a PUD operation, we have a good chance of having a single subarray/mat holding memory regions for the remaining operands of a PUD operation. Fifth, since memory regions might come from different huge pages, we must perform re-mmap to map such memory regions into contiguous virtual addresses.
Mat Label Translation. To keep track of the mapping between mat labels and allocated mat ranges, MIMDRAM adds a small mat translation table alongside the page table. The table is indexed by hashing the mat label with the process ID. It stores in each entry the associated mat range that the memory allocator assigned to that particular mat label. When the CPU dispatches a bbop, the CPU (i) accesses the mat translation table to obtain the mat range assigned to the given bbop, and (ii) replaces the mat label with the mat range.
7 Methodology
We implement MIMDRAM using the gem5 simulator [208] and compare it to a real multicore CPU (Intel Skylake [209]), a real high-end GPU (NVIDIA A100 [210]), and a state-of-the-art PUD framework (SIMDRAM [101]). disable,disable,todo: disable,#D1In all our evaluations, the CPU code is optimized to leverage AVX-512 instructions [157]. Table 2 shows the system parameters we use. To measure CPU energy consumption, we use Intel RAPL [211]. We capture GPU kernel execution time that excludes data initialization/transfer time. To measure GPU energy consumption, we use the nvml API [212]. We implement SIMDRAM on gem5, taking into account that the latency of executing the back-to-back ACTs is only 1.1 the latency of [61, 62, 76, 74, 73, 65], and validate our implementation rigorously with the results reported in [101]. We use CACTI [213] to evaluate MIMDRAM and SIMDRAM energy consumption, where we take into account that each additional simultaneous row activation increases energy consumption by 22% [61, 101]. Our simulation accounts for the additional latency imposes by MIMDRAM’s mat isolation transistors and row decoder latches (i.e., measured (using CACTI [213, 214]) to incur less than 0.5% extra latency for an ACT). We open-source our simulation infrastructure at https://github.com/CMUSAFARI/MIMDRAM.
| Real Intel Skylake CPU [209] | x86 [199], 16 cores, 8-wide, out-of-order, ; |
| L1 Data + Inst. Private Cache: , 8-way, line; | |
| L2 Private Cache: , 4-way, line; | |
| L3 Shared Cache: , 16-way, line; | |
| Main Memory: DDR4-2133, 4 channels, 4 ranks | |
| Real NVIDIA A100 GPU [210] | technology node; 6912 CUDA Cores; |
| 108 streaming multiprocessors, base clock; | |
| L2 Cache: L2 Cache; Main Memory: HBM2 [119, 120] | |
| Simulated SIMDRAM [101] & MIMDRAM | gem5 system emulation; x86 [199], 1-core, out-of-order, ; |
| L1 Data + Inst. Cache: , 8-way, line; | |
| L2 Cache: , 4-way, line; | |
| Memory Controller: row size, FR-FCFS [215, 216] | |
| Main Memory: DDR4-2400, 1 channel, 8 chips, 4 rank | |
| 16 banks/rank, 16 mats/chip, 1 K rows/mat, 512 columns/mat | |
| MIMDRAM’s Setup: 8 entries mat queue, bbop buffer | |
| 8 µProgram processing engines, mat translation table |
Real-World Applications. We analyze 117 applications from seven benchmark suites (SPEC 2017 [164], SPEC 2006 [217], Parboil [218], Phoenix [161], Polybench [162], Rodinia [163], and SPLASH-2 [219]) to select applications that (i) are memory-bound, and (ii) the most time-consuming loop can be auto-vectorized. From this analysis, we collect 12 multi-threaded CPU applications (disable,disable,todo: disable,#D3as Table 3 describes) from different domains (i.e., video compression, data mining, pattern recognition, medical imaging, stencil computation), and their respective GPU implementations, when available. Our evaluated applications are: (i) 525.x264_r (x264) from SPEC 2017; (ii) heartwall (hw), kmeans (km), and backprop (bs) from Rodinia; (iii) pca from Phoenix; and (iv) 2mm, 3mm, covariance (cov), doitgen (dg), fdtd-apml (fdtd), gemm (gmm), and gramschmidt (gm) from Polybench.141414Several prior works [220, 189, 72, 128, 221, 222, 223] show disable,disable,todo: disable,#E1 that our selected twelve workloads can benefit from different types of PIM architectures. Since our base PUD substrate (SIMDRAM) does not support floating-point, we manually modify the selected floating-point-heavy auto-vectorized loops to operate on fixed-point data arrays.151515disable,disable,todo: disable,#D10We only modify the three applications from the Rodinia benchmark suite to use fixed-point operations. Prior works [128, 224, 225] also employ fixed-point for the same three Rodinia applications. The applications from Polybench can be configured to use integers; the auto-vectorized loops in 525.x264_r use uint8_t; pca uses integers. We do not observe an output quality degradation when employing fixed-point for the selected loops. We use the largest input dataset available and execute each application end-to-end in our evaluations.disable,disable,todo: disable,#E1
pca (pca) | reference | 2 | {4000, 4000} | D, S, M, R | ||||||||||||
| Polybench [162] | 2mm (2mm) |
6 | {4000, 4000} | M, R | ||||||||||||
| 3mm (3mm) |
7 | {4000, 4000} | M, R | |||||||||||||
| covariance (cov) |
2 | {4000, 4000} | D, S, R | |||||||||||||
| doitgen (dg) |
5 | {1000, 1000} | M, C, R | |||||||||||||
| fdtd-apml (fdtd) |
3 | {1000, 1000} | D, M, S, A | |||||||||||||
| gemm (gmm) |
4 | {4000, 4000} | M, R | |||||||||||||
| gramschmidt (gs) |
5 | {4000, 4000} | M, D, R | |||||||||||||
| Rodinia [163] | backprop (bs) | 134217729 input elm. | 1 | {17, 134217729} | M, R | ||||||||||||
| heartwall (hw) | reference | 4 | {1, 2601} | M, R | |||||||||||||
| kmeans (km) | 16384 data points | 2 | {16384, 16384} | S, M, R | |||||||||||||
525.x64_r (x264) |
2 | {64, 320} | A |
: D = division, S = subtraction, M = multiplication, A = addition, R = reduction, C = copy
: application with independent PUD operations
Multi-Programmed Application Mixes. To measure system throughput and fairness, we manually create 495 application mixes by randomly selecting eight applications (from our group of 12 applications) for execution co-location. We classify each application mix into one of three categories: low, medium, and high vectorization factor (VF) mixes based on Fig. 3. In the low mix, the maximum VF of all eight applications is lower than 16K; in the medium mix, at least one application has a maximum vectorization factor between 16K (inclusive) and 64K; and in the large mix, at least one application has a maximum VF larger than 64K (inclusive).
Comparison to State-of-the-Art PIM Architectures. We compare MIMDRAM to two other state-of-the-art PIM architectures: DRISA [63] and Fulcrum [104]. DRISA is a combined PUM and PNM architecture that significantly modifies the DRAM microarchitecture and organization to enable bulk in-DRAM computation (e.g., by using 3T1C DRAM cells to execute in-situ bitwise NOR operations and by adding logic gates near the subarray’s sense amplifiers). Fulcrum is a PNM architecture that adds computation logic near subarrays. Fulcrum’s primary components are a series of shift registers (called walkers) that latch input/output DRAM rows and a narrow scalar ALU that executes arithmetic and logic operations. We model the DRISA 3T1C implementation and Fulcrum (i) using a DRAM module of equal dimensions (i.e., number of DRAM ranks, chips, banks, mats, rows, and columns) as the baseline DDR4 DRAM we use for SIMDRAM and MIMDRAM (see Table 2) and (ii) including all the changes that the DRISA 3T1C and Fulcrum architectures propose to the DRAM cell array and DRAM subarray.
8 Evaluation
We demonstrate the advantages of MIMDRAM by evaluating (i) SIMD utilization and energy efficiency (i.e., performance per Watt) for single applications (section 8.1); (ii) system throughput (in terms of weighted speedup [226, 227, 228]), job turnaround time (in terms of harmonic speedup [229, 227]), and fairness (in terms of maximum slowdown [230, 231, 232, 233, 234, 215, 235, 236, 237, 238, 239]) for multi-programmed application mixes in comparison to the baseline CPU, GPU, and a state-of-the-art PUD architecture, i.e., SIMDRAM [101] (section 8.2); (ii) area-normalized performance analysis for single applications and throughput analysis for multi-programmed application mixes in comparison to state-of-the-art PIM architectures, i.e., DRISA [63] and Fulcrum [104] (section 8.3). In most of our analyses (section 8.1–section 8.2), to keep our analyses pure, we very conservatively allow MIMDRAM to use only a single DRAM subarray in a single DRAM bank for PUD computation. In section 8.4, we perform a scalability analysis to evaluate MIMDRAM’s performance when enabling multiple DRAM subarrays and banks for PUD computation, which reflects a more accurate evaluation of the true benefits of MIMDRAM and PUD. Finally, we evaluate MIMDRAM’s DRAM and CPU area cost (section 8.5).
8.1 Single-Application Results
Fig. 9 shows MIMDRAM’s SIMD utilization and normalized energy efficiency (in performance per Watt) for all 12 applications. Values are normalized to the baseline CPU.
SIMD Utilization. We make two observations from Fig. 9a. First, MIMDRAM significantly improves SIMD utilization over SIMDRAM. On average across all applications, MIMDRAM provides 15.6 the SIMD utilization of SIMDRAM. This is because MIMDRAM matches the available SIMD parallelism in an application with the underlying PUD resources (i.e., PUD SIMD lanes) by using only as many DRAM mats as the maximum vectorization factor of a given application’s loop. In contrast, SIMDRAM always occupies all available PUD SIMD lanes (i.e., entire subarrays) for a given operation, resulting in low SIMD utilization for applications without a very-wide vectorization factor. Second, we observe that SIMD utilization can vary considerably within an application. For example, MIMDRAM’s SIMD utilization for hw and bp goes from as low as 0.2% to as high as 100%. This happens because the SIMD parallelism for each vectorized loop in these applications changes at different execution phases. MIMDRAM can better adjust to the variation in SIMD parallelism (than SIMDRAM) due to its flexible design. We conclude that MIMDRAM greatly improves overall SIMD utilization for many applications.
Performance & Energy Efficiency. We make three observations from Fig. 9b. First, MIMDRAM significantly improves energy efficiency and performance over SIMDRAM. On average across all applications, MIMDRAM provides 14.3 the energy efficiency and 34 the performance of SIMDRAM. MIMDRAM’s higher energy efficiency is due to three main reasons. (i) MIMDRAM parallelizes the computation of independent bbops in a single application loop across different mats, improving overall performance. MIMDRAM reduces execution time disable,disable,todo: disable,#C1by 2.8 compared with SIMDRAM, on average across applications with independent bbops (i.e., pca, 3mm, and fdtd). (ii) MIMDRAM implements in-situ PUD vector reduction operations, while SIMDRAM requires the assistance of the CPU for vector reduction, increasing latency and energy consumption. disable,disable,todo: disable,#C1MIMDRAM reduces execution time and energy consumption by 1.6 and 266 over SIMDRAM, on average across the applications with vector reduction operations (from our twelve applications, only fdtd and x264 do not require vector reduction operations). (iii) MIMDRAM activates only the necessary PUD SIMD lanes during an application loop’s execution, significantly saving energy when the application has low SIMD utilization. disable,disable,todo: disable,#C1MIMDRAM reduces energy consumption by 325 over SIMDRAM, on average across applications with a maximum vectorization factor lower than 65,536 (from our twelve applications, only bs exhibits a vectorization factor higher than 65,536). disable,disable,todo: disable,#D2Second, MIMDRAM provides 30.6/6.8 the energy efficiency of CPU/GPU baselines. MIMDRAM’s higher energy efficiency is due to its inherent ability to avoid costly data movement operations for memory-bound applications. Third, even though MIMDRAM improves performance (by 3.1, 8.6, 1.1, and 1.3) compared to the baseline CPU for some applications (i.e., 2mm, cov, gs, and bp), it leads to performance loss compared to the baseline CPU and GPU on average across all applications. This is because, for some applications, the bulk parallelism available inside a single DRAM subarray and bank is insufficient to hide the latency of costly bit-serial operations (e.g., multiplication). We observe that enabling MIMDRAM in 16 DRAM banks and 64 subarrays (per bank) allows MIMDRAM to provide performance gains compared to the CPU and the GPU (see section 8.4). We conclude that MIMDRAM is an energy-efficient and high-performance PUD system.
8.2 Multi-Programmed Workload Results
We evaluate SIMDRAM and MIMDRAM’s impact on system throughput (in terms of weighted speedup [226, 227, 228]), job turnaround time (in terms of harmonic speedup [229, 227]), and fairness (in terms of maximum slowdown [230, 231, 232, 233, 234, 215, 235, 236, 237, 238, 239]) when executing applications concurrently. To provide a fair comparison, we introduce MIMD parallelism in SIMDRAM with bank-level parallelism (BLP) [240, 230, 14, 241, 242], where each SIMDRAM-capable DRAM bank can independently run an application. We evaluate four configurations of SIMDRAM where 1 (SIMDRAM:1), 2 (SIMDRAM:2), 4 (SIMDRAM:4), and 8 (SIMDRAM:8) banks have SIMDRAM computation capability. Fig. 10 shows the system throughput, job turnaround time (which measures a balance of fairness and throughput), and fairness that SIMDRAM and MIMDRAM provide on average across all application mixes. Values are normalized to SIMDRAM:1. We make three observations.
First, MIMDRAM significantly improves system throughput, job turnaround time, and fairness compared with SIMDRAM. On average, across all application groups, MIMDRAM achieves: (i) 1.68 (min. 1.52, max. 2.02) higher weighted speedup, (ii) 1.33 (min. 1.17, max. 1.72) higher harmonic speedup, and (iii) 1.32 (min. 0.95, max. 2.29) lower maximum slowdown than SIMDRAM (averaged across all four configurations). Second, MIMDRAM using a single subarray and single bank for computation, provides 1.68, 1.54, and 1.52 the system throughput of SIMDRAM using 2, 4, and 8 banks for computation, respectively. This happens because MIMDRAM (i) utilizes idle resources at DRAM mat granularity to execute computation as soon as a mat is available, thus reducing queuing time and improving parallelism; and (ii) reduces execution latency of a single application due to its concurrent execution of independent bbop instructions and support for PUD vector reduction. Third, even though MIMDRAM achieves similar fairness compared with SIMDRAM:4 and SIMDRAM:8 for application mixes with low and medium VF, MIMDRAM’s maximum slowdown is 15% (12%) higher than SIMDRAM:8 (SIMDRAM:4) for application mixes with high VF. This is because in MIMDRAM (i) applications share the DRAM mats available inside a single DRAM bank and (ii) bbops are dispatched to execution using an online first fit algorithm. In this way, an application in a mix with high occupancy and execution latency penalizes an application with low occupancy and execution time, negatively impacting fairness. In contrast, such interference does not happen in SIMDRAM:8 since each application is assigned to a different DRAM bank to execute at the cost of occupying eight banks instead of one. MIMDRAM’s fairness can be further improved by (i) employing better scheduling algorithms that target quality-of-service [240, 229, 243, 234, 232, 236] or (ii) using subarray-level parallelism (SALP) [14] and BLP [240, 230, 14, 241, 242] in MIMDRAM, i.e., exploiting multiple subarrays and multiple banks for MIMDRAM computation (section 8.4).161616In our extended version [244], we provide multi-programmed workload results while exploiting SALP and BLP for MIMDRAM computation. We conclude that MIMDRAM is an efficient and high-performance PUD substrate when the system concurrently executes several applications.
CPU Multi-Programmed Workload Results. We evaluate how MIMDRAM performance compares to that of a state-of-the-art CPU when executing multiple applications. To do so, we randomly generate ten different application mixes, each containing eight applications out of our 12 applications. Then, we run each application mix in our baseline CPU (using multi-threading) and in MIMDRAM and compute the achieved system throughput for each system (using weighted speedup). Fig. 11 shows the system throughput MIMDRAM achieves compared to the baseline CPU. We observe that MIMDRAM improves overall throughput by 19%. This is because MIMDRAM can parallelize the execution of the applications in each application mix across the DRAM mats in a subarray. In contrast, when executing each application mix, the baseline CPU often suffers from contention in its shared resources (e.g., shared cache and DRAM bus). We conclude that MIMDRAM is an efficient substrate for highly-parallel environments.
8.3 Comparison to Other PIM Architectures
Single-Application Results. We compare the performance of each PIM architecture and MIMDRAM. Since DRISA and Fulcrum use large additional area (i.e., 21% and 82% DRAM area overhead, respectively, over our baseline DDR4 DRAM chip) to implement PIM operations, we report area-normalized results (i.e., performance per area) for a fair comparison. We use the area values reported in both DRISA and Fulcrum’s papers, scaled to the baseline DDR4 DRAM device we employ. We allow each mechanism to leverage the data parallelism available in each application by dividing the work evenly across DRISA’s PIM-capable DRAM banks and Fulcrum’s PIM-capable subarrays. Fig. 12 shows the normalized performance per area for all 12 applications. Values are normalized to MIMDRAM. We make two observations. First, MIMDRAM achieves the highest performance per area compared to DRISA and Fulcrum. On average across the 12 applications, MIMDRAM performance per area is 1.18/1.92 that of DRISA and Fulcrum. This is because although DRISA and Fulcrum achieve higher absolute performance than MIMDRAM (7.5 and 3.0, respectively), such performance benefits come at the expense of very large area overheads. disable,disable,todo: disable,#CQ1While MIMDRAM incurs small area cost on top of a DRAM array disable,disable,todo: disable,#C1(1.11% DRAM area overhead, see section 8.5), DRISA and Fulcrum incur significantly larger area costs. Second, for some applications (namely hw, dg, km, and x264), DRISA and Fulcrum achieve higher performance per area than MIMDRAM. We observe that such applications are dominated by multiplication operations, which are costly to implement using MIMDRAM’s bit-serial approach. We conclude MIMDRAM is an area-efficient PIM architecture, which provides performance benefits compared to state-of-the-art PIM architectures for a fixed area budget.
Multi-Programmed Workload Results. Fig. 13 shows the system throughput, job turnaround time, and fairness that DRISA, Fulcrum, and MIMDRAM provide on average across all application mixes. We employ BLP in DRISA and MIMDRAM, and SALP [14] in Fulcrum to enable MIMD execution. disable,disable,todo: disable,#CQ1We make two observations. First, all three PIM architectures achieve similar system throughput. On average across all application mixes and configurations, DRISA, Fulcrum, and MIMDRAM achieve 1.20, 1.17, and 1.19 the system throughput of DRISA:1, respectively. Second, when considering a single DRAM subarray for computation, MIMDRAM achieves 8% and 11% higher fairness than DRISA and Fulcrum, respectively.
8.4 MIMDRAM with SALP & BLP
One of the main advantages of PUD architectures is the ability to exploit the large internal DRAM parallelism for computation. A PUD substrate can leverage SALP [14] and BLP [240, 230, 14, 241, 242] techniques to operate simultaneously exploit the many DRAM subarrays (e.g., 8–64 per bank) and banks (e.g., 8–16 per rank) in a DRAM chip for PUD computation. To this end, we perform a sensitivity analysis of SIMDRAM and MIMDRAM’s performance for our twelve applications when using multiple DRAM subarrays (1–64 per bank) and DRAM banks (1–16 per rank) for PUD computation, as Fig. 14 depicts. We make two observations from the figure. First, by fully leveraging the internal DRAM parallelism in a DRAM chip, MIMDRAM can provide significant performance gains compared to the baseline CPU. On average across all twelve applications, MIMDRAM (using 64 DRAM subarrays per bank and 16 banks for PUD computation) achieves 13.2 the performance of the CPU (and 2 the performance of the GPU, not shown in the figure). Second, in contrast, SIMDRAM fails to outperform the baseline CPU, even when fully utilizing the internal DRAM for computation (0.08 the performance of the CPU when using 64 DRAM subarrays per bank and 16 banks). This is because: (i) MIMDRAM unlocks further parallelism by leveraging idle DRAM mats for computation and (ii) MIMDRAM reduces the latency of costly vector reduction operations. Third, we observe that MIMDRAM can lead to performance loss compared to the baseline CPU for some workloads, even when using all available DRAM subarrays and banks for computation, for two main reasons: (i) quadratically-scaling PUD operations (i.e., multiplication and divisions) or (ii) PUD vector reduction operations dominate MIMDRAM’s execution time of the application. In the first case, MIMDRAM’s performance could be further improved by leveraging lower-latency algorithms for costly PUD operations (e.g., bit-parallel multiplication and division algorithms [245]) or performing such complex operations near memory (close to DRAM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 104]. In the second case, MIMDRAM would benefit from the assistance of PNM architectures to perform faster vector reduction operations in DRAM, at the cost of an increase in area cost. We conclude that MIMDRAM highly benefits from exploiting SALP and BLP for PUD computation.
8.5 Area Analysis
We use CACTI [213, 214] to model the area of a DRAM chip (Table 2) using a technology node. We implement MIMDRAM’s chip select and mat identifier logic in Verilog HDL and synthesize the HDL using the Synopsys Design Compiler [246] with a process technology node.171717disable,disable,todo: disable,#B1We use a technology node since that is the best CMOS standard cell library we have access to in our environment. We scaled our design to a technology node following prior works’ methodology [247, 24, 248, 249].
DRAM Bank Area. We evaluate the area overhead of (i) mat isolation transistors, (ii) row decoder latches, (iii) mat selectors, (iv) the wiring to propagate mat selector output to mat isolation transistors (matlines), and (v) multiplexers and wiring of the inter-mat interconnect. MIMDRAM incurs 1.15% area overhead over the baseline DRAM bank.
DRAM Chip I/O Area. The total area overhead for MIMDRAM’s chip select and mat identifier is only at a technology node. We estimate the equivalent area overhead at a technology node to be [247].
Overall, MIMDRAM increases the area of the evaluated DRAM chip (16 banks and I/O) by only 1.11%.
MIMDRAM Control Unit & Transposition Unit Area. The main components in the MIMDRAM control unit are the (i) bbop buffer, (ii) mat scoreboard, and (iii) µProgram processing engines. We set the size of the bbop buffer to , which accommodates up to 1024 bbops. The mat scoreboard requires 128 bits of storage, one bit per DRAM mat per subarray. A single µProgram processing engine has an area of . We empirically include eight µProgram processing engines in our design. We estimate, using CACTI, that MIMDRAM control unit area is . MIMDRAM transposition unit has an area equal to the SIMDRAM transposition unit (of ). Considering the area of the control and transposition units, MIMDRAM has a low area overhead of 0.6% over the die area of a state-of-the-art Intel Xeon E5-2697 v3 CPU [72].
9 Related Work
To our knowledge, MIMDRAM is the first end-to-end processing-using-DRAM (PUD) system for general-purpose applications that executes operations in a multiple-instruction multiple-data (MIMD) fashion, where independent PUD operations are executed concurrently across the DRAM mats of a DRAM subarray. We highlight MIMDRAM’s key contributions by contrasting it with state-of-the-art processing-in-memory (PIM) designs. We already compared MIMDRAM to SIMDRAM [101], DRISA [63], and Fulcrum [104] both quantitatively and qualitatively in section 8 and demonstrated MIMDRAM’s benefits over them.
Processing-Using-DRAM. Prior works propose different ways of implementing PUD operations, either by (i) using the memory arrays themselves to perform operations in bulk [74, 64, 75, 182, 61, 67, 52, 66, 69, 63, 101, 88, 78, 79, 145, 83, 85] or (ii) modifying the DRAM sense amplifier design with logic gates for computation [63, 250]. Since prior PUD architectures execute PUD operations at a coarse granularity (i.e., at the granularity of a DRAM row access), they can suffer from the underutilization issue we highlight. As in MIMDRAM, prior PUD architectures can employ fine-grained DRAM for PUD operations to mitigate underutilization. We believe that the principles employed in MIMDRAM can benefit other PUD architectures, leading to performance, energy-efficiency, and programmability improvements for the underlying PUD substrate.
Programming Support for PUM. Prior works propose programming models for different types of PUM architectures, as (i) CUDA/OpenAcc [251, 252] for in-cache computing [72]; (ii) tensor dataflow graphs for in-ReRAM computing [128]. By enabling fine-grained DRAM, we believe such programming models can be now easily ported to PUD computing (for example, by assuming that each DRAM mat executes a different CUDA thread block). CHOPPER [105] improves SIMDRAM’s programming model by leveraging bit-slicing compilers and employing optimizations to reduce the latency of a µProgram. Even though CHOPPER simplifies programmability compared to SIMDRAM, it still requires the programmer to re-write applications using the bit-slicing compiler’s syntax. Compared to CHOPPER, MIMDRAM has two main advantages. First, MIMDRAM automatically generates code for the PUD engine without any code refactoring. Second, since CHOPPER maintains the very-wide SIMD programming model of SIMDRAM, it also suffers from SIMD underutilization.
10 Conclusion
We introduce MIMDRAM, a hardware/software co-designed processing-using-DRAM (PUD) substrate that can allocate and control only the needed computing resources inside DRAM for PUD computing. On the hardware side, MIMDRAM introduces simple modifications to the DRAM architecture that enables the execution of (i) different PUD operations concurrently inside a single DRAM subarray in a multiple-instruction multiple-data (MIMD) fashion, and (ii) native vector reduction computation. On the software side, MIMDRAM implements a series of compiler passes that automatically identify and map code regions to the underlying PUD substrate. We experimentally demonstrate that MIMDRAM provides significant benefits over state-of-the-art CPU, GPU, and processing-using-memory (PUM) and processing-near-memory (PNM) systems. We hope and believe that our work can inspire more efficient and easy-to-program PUD systems. The source code of MIMDRAM is freely available at https://github.com/CMU-SAFARI/MIMDRAM.
Acknowledgments
We thank the anonymous reviewers of MICRO 2023 and HPCA 2024 for their encouraging feedback. We thank the SAFARI Research Group members for providing a stimulating intellectual environment. We acknowledge the generous gifts from our industrial partners; including Google, Huawei, Intel, Microsoft. This work is supported in part by the Semiconductor Research Corporation, the ETH Future Computing Laboratory, the AI Chip Center for Emerging Smart Systems (ACCESS), and the UIUC/ZJU HYBRID Center.

### 03 Mar 2024 v2

@@
 Data movement between computation units (e.g., CPUs, GPUs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], and is expected to worsen due to the increasing data intensiveness of modern applications [21, 22].
-To mitigate the overheads caused by data movement, several works propose processing-in-memory (PIM) architectures [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117].
-There are two main approaches to PIM [8, 9]: (i) processing-near-memory (PNM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 17, 10, 38, 39, 40, 41, 42, 18, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 104], where computation logic is added near the memory arrays (e.g., in a DRAM chip or at the logic layer of a 3D-stacked memory [118, 119, 120]); and (ii) processing-using-memory (PUM) [59, 60, 61, 62, 63, 64, 65, 66, 67, 121, 68, 69, 70, 71, 72, 52, 73, 74, 75, 76, 77, 78, 79, 122, 123, 80, 81, 82, 83, 124, 84, 85, 86, 87, 125, 88, 126, 127, 128, 89, 90, 129, 91, 92, 93, 94, 95, 130, 131, 132, 133, 134, 135, 136, 137, 138, 96, 139, 140, 97, 141, 142, 143, 98, 99, 144, 100, 101, 145, 102, 103, 105, 146], where computation is performed by exploiting the analog operational properties of the memory circuitry.
+To mitigate the overheads caused by data movement, several works propose processing-in-memory (PIM) architectures [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 17, 49, 10, 50, 51, 52, 53, 54, 55, 56, 18, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190].
+There are two main approaches to PIM [8, 9]: (i) processing-near-memory (PNM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 17, 49, 10, 50, 51, 52, 53, 54, 55, 56, 18, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 176], where computation logic is added near the memory arrays (e.g., in a DRAM chip or at the logic layer of a 3D-stacked memory [191, 192, 193]); and (ii) processing-using-memory (PUM) [106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 73, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 177, 183], where computation is performed by exploiting the analog operational properties of the memory circuitry.
 There are two main advantages of PUM over PNM.
@@
 Second, PUM architectures exploit the large internal bandwidth and parallelism available inside the memory arrays, while PNM solutions are fundamentally bottlenecked by the memory’s internal data buses.
-PUM architectures can be implemented using different memory technologies, including SRAM [71, 70, 96, 97, 98, 99], DRAM [125, 52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], emerging non-volatile [77, 60, 68, 82, 81, 90, 59] or NAND flash [80, 91, 92, 93, 94, 95].
-Processing-using-DRAM (PUD) [52, 85, 66, 78, 79, 69, 63, 88, 83, 86, 74, 64, 76, 73, 65, 61, 75, 62, 67], in particular, enables the execution of different bulk operations in DRAM (i.e., PUD operations), such as (i) data copy and initialization [74, 64, 75, 100], (ii) bitwise Boolean operations [61, 69, 67, 52, 63], (iii) arithmetic operations [66, 69, 63, 125, 101, 88], and (iv) lookup table based operations [78, 79, 145, 102, 103].
+PUM architectures can be implemented using different memory technologies, including SRAM [119, 118, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169], DRAM [139, 73, 136, 113, 126, 127, 117, 110, 140, 133, 137, 122, 111, 124, 121, 112, 108, 123, 109, 114], emerging non-volatile [128, 125, 107, 116, 115, 132, 131, 145, 106, 170] or NAND flash [130, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158].
+Processing-using-DRAM (PUD) [139, 73, 136, 113, 126, 127, 117, 110, 140, 133, 137, 122, 111, 124, 121, 112, 108, 123, 109, 114], in particular, enables the execution of different bulk operations in DRAM (i.e., PUD operations), such as (i) data copy and initialization [122, 111, 123, 171], (ii) bitwise Boolean operations [108, 117, 114, 73, 110], (iii) arithmetic operations [113, 117, 110, 139, 172, 140], and (iv) lookup table based operations [126, 127, 173, 174, 175].
 PUD architectures commonly employ bit-serial computation [139, 73, 117, 122, 111, 124, 121, 112, 108, 123, 109, 114, 172], where they map each data element of a PUD operation to a DRAM column.111We provide a detailed background on DRAM organization in section 2.1.